OpenStudy (anonymous):

Solve the matrix equation for a, b, c, and d. (a-b) (b+a) (3d+c) (2d-c) Equals 8 1 7 6

OpenStudy (bahrom7893):

a-b = 8 b+a = 1 So, a = 8+b, b+8+b=1, 2b = -7, b = -7/2

OpenStudy (bahrom7893):

a = 8-7/2 = 9/2

OpenStudy (bahrom7893):

repeat for 3d+c and 2d-c